for equation y-3=3(x+1), i need to write an equation in standard form of a parallel line that passes thru (0,-2) and on the parallel line, i need to find the ordered pair where x=-2. please help
Answer by rothauserc(4718)

the standard form for a line is y = mx + b where m is the slope and b is the y intercept
y-3=3(x+1)
y = 3x + 3 +3
y = 3x + 6
slope of the line is 3
the parallel line will have the same slope
y = 3x + b
we solve for b by using the point (0, -2)
-2 = 3*0 + b
b = -2 and the formula for the parallel line is
y = 3x -2
now for the ordered pair when x = -2
y = 3*-2 -2
y = -8
the ordered pair is (-2, -8)